What is the adjusted body weight for a woman who is 5’4” and 170#?

What is the adjusted body weight for a large-framed man who is 72” and 285#?
Answer #1

Adjusted body weight.

Adjusted body weight is used commonly in clinical practice to determine doses of the drugs, absorption, metabolism, excretion, and distribution of the drugs, volume adjustment in a patient with kidney problems, in mild to the moderately obese patient.

GIVEN DATA ARE.

WEIGHT OF THE PATIENT ( actual)= 170 KGS.

HEIGHT OF THE PATIENT = 5" 4

= 162.56 cms.

~ 163 cm

To calculate adjusted body weight we need to calculate the ideal body weight first.

IDEAL BODY WEIGHT= 45.5 ( for females) + 2.3 ( with each inches above 5 foot)

As patients height is. 5" 4, there are 4 inches that are to be

Multiplied with 2.3.

So, 4 * 2.3= 9.2 kgs

IBW ( ideal body weight ) = 45.5 kgs + 9.2 kgs

= 54.7. ~ 55 kgs.

The formula for adjusted body weight.

Adbw= IBW +0.4 ( actual body weight - IBW)

= 55 +0.4 ( 170 - 55 )

= 55 + 0.4 ( 115)

= 55 + 46.

= 101 kgs.

Adjusted body weight will be 101 kgs.

For men

Given data are

Weight ( actual body weight) = 285 kgs.

Height = 72"

= 6 feet.

Ideal body weight IBW= 50 + 2.3 ( over each inches above 5 feet).

The patient is 6 feet, so 1 foot will be multiplied with 2.3.

1 feet = 12 inches.

= 12* 2.3. = 27.6

IBW= 50+ 27.6

= 77.6 kgs.

Adjusted body weight = IBW+ 0.4 ( Actual body weight - IBW)

= 77.6 +0.4 ( 285 - 77.6)

= 77.6 + 0.4 ( 207.4)

= 77.6 + 82.96

= 160.56 kgs

Actual body weigh ~ 161 kgs.